how do u solve (-3x)-y=9 in slope intercept form?
Answer by richwmiller(17219)

Slope intercept form is y=mx+b
-3x-y=9
becomes
-x-9=y
y=-3x-9
where m is the slope and b is the y intercept
You can get the slope in the form it is in now.
-3x-y=9
a=-3
b=-1
note: this b is not the same as the b of y=mx+b
m=-a/b=-(-3)/-1=-3
That was fun with three minuses.